OpenStudy (anonymous):

Equation M: y = x + 4 Equation P: y = 3x + 6 Which of the following is a step that can be used to find the solution to the set of equations? A. x=3x+6 B. x+4=3x+6 C. x+6=3x+4 D. x=3x HELPPPPPPPP!!!!!!!!!!!!!!

OpenStudy (anonymous):

@Godlovesme @LazyBoy

OpenStudy (anonymous):

@jeffyblood

OpenStudy (anonymous):

@Nnesha

Directrix (directrix):

y = x + 4 and y = 3x + 6 If both x + 4 and 3x + 6 are equal to y, what does that say about x + 4 and 3x + 6 @neee

OpenStudy (anonymous):

there equal too

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Godlovesme

Directrix (directrix):

>> there equal too Correct So, x+4=3x+6 @neee

OpenStudy (anonymous):

thank you